# Member, IEEE

### v1

@@ Abstract
{IEEEkeywords}
Two-Body, Circular Restricted Three-Body, Koopman Theory, Neural Network, Linearization
@@ 1 INTRODUCTION
The rapid growth in motivation to explore space in the last fifteen years has sparked a number of significant missions to planets such as Mars and Venus, moons such as Enceladus and Titan, as well as our own Moon. From interplanetary missions to the placement of satellites within our own planet’s orbit, we would not be able to achieve the goals of these endeavors without understanding and exploiting the natural dynamics of these systems. Much effort has been spent in determining how exactly the dynamics of bodies, either natural or man-made, are influenced, and how engineers can develop guidance, navigation, and control systems to complete these extraordinary tasks. Today, thousands of satellites orbit around our Earth in a variety of different orbits, and the governing equations of motion for all of these satellites are the equations originating from the N-Body Problem. In the N-Body equations of motion, the only force that imparts energy onto the bodies is gravity [1], and dependent on the complexity of the solution required and the nature of the system in question, the number of bodies varies. Most often in celestial dynamics, the Two-Body Problem (2BP) and Circular Restricted Three-Body Problem (CR3BP) are analysed and used [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As we endeavour to reach the Moon again, and move forth to Mars, the analysis and use of the CR3BP will become even more prevalent [7, 10, 11].
@@
Koopman theory, first proposed in 1931 [17], has gained traction over the last few years as a solution for finding global linearization of nonlinear systems. In a nutshell, the theory states that the dynamics of a nonlinear system can be described linearly by an infinite-dimensional Koopman operator. Due to its infinite dimension, for practical use, it is typically approximated using data-driven methods such as Extended Dynamic Mode Decomposition (EDMD) [18]. The process of implementing EDMD on the nonlinear system to find an accurate approximation is not only highly involved, but requires an innate understanding of the method in order to select the basis functions. These basis functions are the building blocks from which functions known as observable functions are constructed. These observable functions are then what is applied to the data in order to lift dynamics to the appropriate dimension. However, constructing a feed-forward neural network (NN) that implements EDMD makes it possible to find an approximate general Koopman operator linearization applicable to a region of state space, that is time invariant and constant. Another advantage of using a data-driven method is that it does not require system knowledge, and the nonlinear system can be completely unknown [19, 20].
@@ 2.1.3 Koopman Approximation Algorithm
[…]
@@ 2.2 Circular Orbital Dynamics
$\mathbf{x}=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\\ x_{4}\end{bmatrix}=\begin{bmatrix}x\\ y\\ \dot{x}\\ \dot{y}\end{bmatrix}$ (20)
@@
$\mathbf{x_{0}}=\begin{bmatrix}r\\ 0\\ 0\\ \sqrt{\frac{\mu}{r}}\end{bmatrix}$ (23)
@@ 2.3 Circular Restricted Three-Body Dynamics
$\mathbf{x}=\begin{bmatrix}x\\ y\\ z\\ \dot{x}\\ \dot{y}\\ \dot{z}\end{bmatrix}$
@@
$0=$ $-\frac{1-\mu}{(x+\mu)\cdot\left|x+\mu\right|}$ (28)
$-\frac{\mu}{(x-1+\mu)\cdot\left|x-1+\mu\right|}+x$
@@ 4.1 Two-Body Circular Metrics
In addition, the Kepler problem is endowed with invariants such as the Laplace–Runge–Lenz vector that will not be addressed in this work [44].
@@ 5.2.1 Two-Body Problem Linearization
Lifted Space Size
6
Hidden Layers
3
Neurons per hidden layer
25
Batch Size
128
Learning Rate
0.0001
Optimizer
Adam
Activation Function
SELU
Weight Decay
0.00001
Epochs
80000
$\gamma$
0.8
$\beta$
1
$\lambda_{L_{1}}$
0.04
$\lambda_{L_{2}}$
0.01
K matrix dimension
10 x 10
@@ 5.2.3 Circular Restricted Three-Body Problem Linearization
Lifted Space Size
100
Hidden Layers
13
Neurons per hidden layer
105
Batch Size
16
Learning Rate
0.000001
Optimizer
Adam
Activation Function
SELU
Weight Decay
0.00001
Epochs
35000
$\gamma$
2
$\beta$
1
$\lambda_{L_{1}}$
0.004
$\lambda_{L_{2}}$
0.001
K matrix dimension
106 x 106
@@
Figure 12 shows the linearized evolution of the dynamics that is generated by our Koopman operator. It shows that the DNN is able to learn appropriate observables and lift the dimension of the states such that the K matrix was able to accurately predict the orbit of the CR3BP by rendering it as a linear system. Because the orbit, like the 2BP, is in-plane, the $z$-axis position and velocity are zero, therefore we are able to plot the propagation of the dynamics on a 2D plane to further inspect the accuracy of the propagation, as done in Figure 13. Note that in both Figures 12 and 13 the scale of the Earth and Moon representations are chosen for ease of viewing, and are not to scale with the propagated orbit. We also show the evolution of each of the 4 states independently in Figure 14, comparing them to the nonlinear dynamics, to emphasise the ability to closely follow the intricate dynamics of the CR3BP. It shall be noted that although the CR3BP is propagated in the non-dimensionalized units, in these figures the dimensionalized units are restored, and are displayed in the rotating frame.
@@
$\mathbf{x_{0}}=\begin{bmatrix}0.8673\\ 0\\ 0\\ 0\\ -0.2546\\ 0\end{bmatrix}$

### v2

# Deep Learning Based Dynamics Identification and Linearization of Orbital Problems using Koopman Theory

@@ Abstract
[…]
organization=Department of Aerospace, Physics and Space Sciences, Florida Institute of Technology,addressline=150 W University Blvd, city=Melbourne, postcode=32901, state=FL, country=USA
@@ 1 INTRODUCTION
The rapid growth in motivation to explore space in the last fifteen years has sparked a number of significant missions to planets such as Mars and Venus, moons such as Enceladus and Titan, as well as our own Moon. From interplanetary missions to the placement of satellites within our own planet’s orbit, we would not be able to achieve the goals of these endeavors without understanding and exploiting the natural dynamics of these systems. Much effort has been spent in determining how exactly the dynamics of bodies, either natural or man-made, are influenced, and how engineers can develop guidance, navigation, and control systems to complete these extraordinary tasks. Today, thousands of satellites orbit around our Earth in a variety of different orbits, and the governing equations of motion for all of these satellites are the equations originating from the N-Body Problem. In the N-Body equations of motion, the only force that imparts energy onto the bodies is gravity [1], and dependent on the complexity of the solution required and the nature of the system in question, the number of bodies varies. Most often in celestial dynamics, the Two-Body Problem (2BP) and Circular Restricted Three-Body Problem (CR3BP) are analysed and used [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As we endeavour to reach the Moon again, and move forth to Mars, the analysis and use of the CR3BP will become even more prevalent [7, 10, 11, 12, 13].
@@
Koopman theory, first proposed in 1931 [19], has gained traction over the last few years as a solution for finding global linearization of nonlinear systems. In a nutshell, the theory states that the dynamics of a nonlinear system can be described linearly by an infinite-dimensional Koopman operator. Due to its infinite dimension, for practical use, it is typically approximated using data-driven methods such as Extended Dynamic Mode Decomposition (EDMD) [20]. The process of implementing EDMD on the nonlinear system to find an accurate approximation is not only highly involved, but requires an innate understanding of the method in order to select the basis functions. These basis functions are the building blocks from which functions known as observable functions are constructed. These observable functions are then what is applied to the data in order to lift dynamics to the appropriate dimension. However, constructing a feed-forward neural network (NN) that implements EDMD makes it possible to find an approximate general Koopman operator linearization applicable to a region of state space, that is time invariant and constant. Because of the high interest in Koopman Theory as a method for globally linearizing the nonlinear dynamics of a problem, a great deal of research has been conducted in a wide variety of fields. Applications in fluid dynamics [21], robotics [22] and control systems [23] have all been investigated. Whilst the utilization of Koopman in order to develop linear systems capable of control has been researched heavily, with many models including optimal controls [24], Model Predictive Controllers (MPC) [25], Linear Quadratic Controllers (LQR) [26], all being successfully developed through the implementation of EDMD with control (EDMDc) [27] via analytical solutions or the development of NN’s. Another advantage of using a data-driven method is that it does not require system knowledge, and the nonlinear system can be completely unknown [28, 29].
@@
Current state of the art of the linearization of the TBP and CR3BP includes the use of Koopman Operator Theory. The relative, Clohessy-Wiltshire equations were successfully modelled by Koopman Theory to relieve some of the underlying assumptions as shown by [37]. Other models, including an oblate planet problem [38], Sun-synchronous orbit [39] and Earth-Moon L1 Lagrange point system [40] have all been successfully linearized with the Koopman operator. On the other hand, [41, 2, 42] were able to utilize the Koopman operator to study and anaylze the separate challenges related to modelling the CR3BP, including models around colinear Lagrange points and uncertainties in the propagation of the dynamics. The work [43] however, implements EDMD to analyse the CR3BP in regions outside of these Lagrange points, expanding on the breadth of knowledge base of Koopman applications on orbital problems. Because of the numerical complexity of both the TBP and CR3BP, the treatment of the theory in this domain is often contextualised to specific conditions rather than generalized to the entire problem as a whole.
@@ 2.1.3 Koopman Approximation Algorithm
1:Input: X,Y training data, batch size $b_{s},$ Epochmax
@@ 2.2 Circular Orbital Dynamics
$\mathbf{x}=\begin{bmatrix}x_{1}&x_{2}&x_{3}&x_{4}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}x&y&\dot{x}&\dot{y}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ (20)
@@
$\mathbf{x_{0}}=\begin{bmatrix}r&0&0&\sqrt{\frac{\mu}{r}}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ (23)
@@ 2.3 Circular Restricted Three-Body Dynamics
$\mathbf{x}=\begin{bmatrix}x&y&z&\dot{x}&\dot{y}&\dot{z}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$
@@
$0=$ $-\frac{1-\mu}{(x+\mu)\cdot\left|x+\mu\right|}-\frac{\mu}{(x-1+\mu)\cdot\left|x-1+\mu\right|}+x$ (28)
@@ 4.1 Two-Body Circular Metrics
In addition, the Kepler problem is endowed with invariants such as the
Laplace–Runge–Lenz vector that will not be addressed in this work [60].
@@ 5.2.1 Two-Body Problem Linearization
Lifted Space Size 6
Hidden Layers 3
Neurons per hidden layer 25
Batch Size 128
Learning Rate 0.0001
Optimizer Adam
Activation Function SELU
Weight Decay 0.00001
Epochs 80000
$\gamma$ 0.8
$\beta$ 1
$\lambda_{L_{1}}$ 0.04
$\lambda_{L_{2}}$ 0.01
K matrix dimension 10 x 10
@@ 5.2.3 Circular Restricted Three-Body Problem Linearization
Lifted Space Size 100
Hidden Layers 13
Neurons per hidden layer 105
Batch Size 16
Learning Rate 0.000001
Optimizer Adam
Activation Function SELU
Weight Decay 0.00001
Epochs 35000
$\gamma$ 2
$\beta$ 1
$\lambda_{L_{1}}$ 0.004
$\lambda_{L_{2}}$ 0.001
K matrix dimension 106 x 106
@@
Figure 13 shows the linearized evolution of the dynamics that is generated by our Koopman operator. It shows that the DNN is able to learn appropriate observables and lift the dimension of the states such that the K matrix was able to accurately predict the orbit of the CR3BP by rendering it as a linear system. Because the orbit, like the 2BP, is in-plane, the $z$-axis position and velocity are zero, therefore we are able to plot the propagation of the dynamics on a 2D plane to further inspect the accuracy of the propagation, as done in Figure 14. Note that in both Figures 13 and 14 the scale of the Earth and Moon representations are chosen for ease of viewing, and are not to scale with the propagated orbit. We also show the evolution of each of the 4 states independently in Figure 12, comparing them to the nonlinear dynamics, to emphasise the ability to closely follow the intricate dynamics of the CR3BP. It shall be noted that although the CR3BP is propagated in the non-dimensionalized units, in these figures the dimensionalized units are restored, and are displayed in the rotating frame.
@@
$\mathbf{x_{0}}=\begin{bmatrix}0.8673&0&0&0&-0.2546&0\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$